OpenStudy (anonymous):

Let f(x) = -4x + 7 and g(x) = 10x - 6. Find f(g(x)).

Nnesha (nnesha):

\[\huge\rm f(\color{red}{g(x)})\textrm {is same as }f(\color{reD}{10x-6})\] so substitute x for 10x-6 into f(x) function

OpenStudy (anonymous):

so it would be x = 3/5?

Nnesha (nnesha):

no f(g(x)) means you should substitute x for g(x) function which is 10x-6 so in other words replace x by 10x-6 x=10x-6

OpenStudy (anonymous):

ohh okay

Nnesha (nnesha):

\[\huge\rm f(\color{reD}{10x-6})=-4\color{reD}{x}+7\] replace the x by 10x-6

OpenStudy (anonymous):

wait i'm still confused

Nnesha (nnesha):

okay say it:=)

Nnesha (nnesha):

so what's ur question ?

OpenStudy (anonymous):

i don't get what you do after replacing the x with 10x-6

Nnesha (nnesha):

ohh distribute parentheses by 6

Nnesha (nnesha):

\[\huge\rm f(\color{reD}{10x-6})=-4\color{reD}{(10x-6)}+7\] distribute parentheses by 4 and then combine `like` terms

OpenStudy (anonymous):

so then f = -4(20x - 12) + 7 ??

Nnesha (nnesha):

hmm no

Nnesha (nnesha):

multiply both terms in the parentheses by outside number \[\huge\rm a(b+c)=a \times b+a \times c\] like this

Nnesha (nnesha):